The function H described by H(x)=2.75x+71.48 can be used to predict the height, in centimeters, of a woman whose humerus (the bone from the elbow to the shoulder) is x centimeters long.
Predict the height of the woman whose humerus is 35 cm long.

Answer by checkley77(12844)

H(x)=2.75x+71.48 X=35
F(35)=2.75*35+71.45
F(35)=96.25+71.45
F(35)=167.7 CM IS THE PREDICTED HEIGHT OF THE WOMAN WITH A HUMERUS OF 35 CM.